# Physics-Informed Singular-Value Learning for Cross-Covariances Forecasting in Financial Markets

### v1

@@ 4 Discussion and Conclusion
Across controlled synthetic benchmarks and temporally shuffled equity returns, the proposed estimator matches the performance of the optimal rotationally invariant shrinkage, indicating that it has learned a stationary random-matrix-consistent spectral denoiser rather than introducing spurious structure. In the presence of a pronounced market mode, it also improves upon the corresponding analytic shrinkage method, which is consistent with a mode-aware generalization.
On original (unshuffled) market data, the two cleaners remain comparable for moderate universes, but their behavior diverges as $n$ grows. The analytical BBP shrinkage deteriorates sharply with system size and can even underperform the unfiltered MLE, consistent with sensitivity to macroscopic components (notably a market mode) that violate the boundedness conditions underlying the asymptotics. In contrast, the neural estimator remains stable and achieves lower OOS reconstruction error, including for $n>500$ (an out-of-distribution regime relative to training). Removing the market mode substantially reduces the BBP collapse, yet a residual gap to the neural estimator persists in the chronological evaluation and largely disappears under temporal shuffling; taken together, these controls support the view that the architecture preserves the BBP symmetry constraints while using OOS supervision to retain only cross-dependencies that persist across adjacent windows, i.e., a symmetry-preserving singular-value map that interpolates between denoising and a mild predictive correction.

### v2

@@ 4 Discussion and Conclusion
Across controlled synthetic benchmarks and temporally shuffled equity returns, the proposed estimator achieves near-oracle performances, indicating that it has learned a stationary random-matrix-consistent spectral denoiser rather than introducing spurious structure. In settings with a pronounced market mode, it systematically outperforms the analytical BBP shrinkage, consistent with the fact that macroscopic common components violate the boundedness conditions underlying the asymptotic BBP formula and that the learned mapping can adapt to these mode-driven distortions.
On original (unshuffled) market data, the two cleaners remain comparable for moderate universes, but their behavior diverges as $n$ grows. Also in this case, the analytical BBP shrinkage deteriorates with system size and can even underperform the unfiltered MLE. In contrast, the neural estimator remains stable and achieves lower OOS reconstruction error, including for $n>500$ (an out-of-distribution regime relative to training). Removing the market mode substantially reduces the BBP collapse, yet a residual gap to the neural estimator persists in the chronological evaluation and largely disappears under temporal shuffling; taken together, these controls support the view that the architecture preserves the BBP symmetry constraints while using OOS supervision to retain only cross-dependencies that persist across adjacent windows, i.e., a symmetry-preserving singular-value map that interpolates between denoising and a mild predictive correction.